# Hierarchical Clustering With Confidence

### v1

@@ A.1.2 Computing the p-value
Appendix B Proofs
B.1 Proof of Lemma 1
Proof.
Since $D^{(t)}(X)$ is a function of $X$, we have that
$\mathbb{P}(M^{(t)}=M|X=X_{o})=\mathbb{P}(M^{(t)}=M|X=X_{o},D^{(t)}(X)=D^{(t)}(X_{o})).$
Then, note that
$p^{(t)}(M;X_{o})$ $=\mathbb{P}(M^{(t)}=M|X=X_{o})$
$=\mathbb{P}(M^{(t)}=M|X=X_{o},D^{(t)}(X)=D^{(t)}(X_{o}))$
$=\mathbb{P}(M^{(t)}=M|D^{(t)}(X)=D^{(t)}(X_{o})),$
where the last display follows from the definition of $p^{(t)}(M;X_{o})$ in (1), i.e., $p^{(t)}(M;X_{o})$ depends on data $X$ only through the dissimilarities $D^{(t)}(X)$. ∎
B.2 Proof of Proposition 1
Proof.
For $(M,M_{o}^{*})\in\mathcal{M}^{(t)}$ with $M\notin\mathcal{M}^{*}_{o}$ and $M_{o}^{*}\in\mathcal{M}^{*}_{o}$, we have, by definition, that
$d(M_{o}^{*};X_{o})<d(M;X_{o}).$
Now, observe the following:
$\mathbb{P}(M^{(t)}=M|X=X_{o})$ $=\dfrac{\exp\left(-\dfrac{d(M;X_{o})}{\tau^{(t)}}\right)}{\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}}\exp\left(-\dfrac{d(M^{\prime};X_{o})}{\tau^{(t)}}\right)}\times\dfrac{\exp\left(\dfrac{d(M^{*}_{o};X_{o})}{\tau^{(t)}}\right)}{\exp\left(\dfrac{d(M^{*}_{o};X_{o})}{\tau^{(t)}}\right)}$
$=\dfrac{\exp\left(-\dfrac{1}{\tau^{(t)}}(d(M;X_{o})-d(M^{*}_{o};X_{o}))\right)}{\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}}\exp\left(-\dfrac{1}{\tau^{(t)}}(d(M^{\prime};X_{o})-d(X_{o};M^{*}_{o}))\right)}$
$=\frac{\exp\left(-\dfrac{|\mathcal{M}^{(t)}|}{\tau}\times\dfrac{(d(M;X_{o})-d(M^{*}_{o};X_{o}))}{\sum_{M^{\prime}\in\mathcal{M}^{(t)}}d(M^{\prime};X_{o})}\right)}{1+\displaystyle\sum_{M^{\prime}\in\mathcal{M}^{(t)}\setminus M^{*}_{o}}\exp\left(-\dfrac{|\mathcal{M}^{(t)}|}{\tau}\times\dfrac{(d(M;X_{o})-d(M^{*}_{o};X_{o}))}{\sum_{M^{\prime}\in\mathcal{M}^{(t)}}d(M^{\prime};X_{o})}\right)}$
$<\exp\left(-\dfrac{|\mathcal{M}^{(t)}|}{\tau}\times\dfrac{(d(M;X_{o})-d(M^{*}_{o};X_{o}))}{\sum_{M^{\prime}\in\mathcal{M}^{(t)}}d(M^{\prime};X_{o})}\right).$
As $\tau\rightarrow 0$, the right-hand side of the above inequality tends to 0 since $d(M;X_{o})-d(M_{o}^{*};X_{o})>0$. Hence, $p^{(t)}(M;X_{o})\rightarrow 0$ for all $M\notin\mathcal{M}^{*}_{o}$. Equivalently,
$\lim_{\tau\rightarrow 0}\mathbb{P}(M^{(t)}\in\mathcal{M}^{*}_{o}|X=X_{o})=1.$
B.3 Proof of Lemma 2
Proof.
Note that the observed data matrix $X_{o}$ can be decomposed as follows:
$X_{o}$ $=B^{(t)}_{o}X_{o}+W^{(t)}_{o}X_{o}+(I_{n}-B^{(t)}_{o}-W^{(t)}_{o})X_{o}$
The right-hand side of the above display further equals
B.4 Proof of Lemma 3
Proof.
Moreover, under the null hypothesis, we have $\mathbb{E}[B^{(t)}_{o}X]=\mathbb{E}[W^{(t)}_{o}X]=0$. By the properties of the Gaussian distribution with zero mean and isotropic covariance, it follows that $(\|B^{(t)}_{o}X\|_{F},\|W^{(t)}_{o}X\|_{F})\perp\!\!\!\perp\left(\eta^{(t)},\gamma^{(t)}\right)$. Therefore, $R^{(t)}$ as a statistic that is a measurable function of $\|B^{(t)}_{o}X\|_{F}$$\|W^{(t)}_{o}X\|_{F}$ remains independent of $\left(\eta^{(t)},\gamma^{(t)}\right)$.
$R^{(t)}\Big\lvert\mathcal{A}^{(t)}=\mathcal{A}_{o}^{(t)}$
B.5 Proof of Theorem 2
Proof.
Fixing some notation, let
Then, by the probability integral transform, we obtain that
$\mathbb{P}_{\overline{M}_{o}^{(t)}}(\mathcal{A}^{(t)}_{o})=\alpha.$
Finally, we apply the law of iterated expectation to conclude that
B.6 Proof of Theorem 3
Lemma 6.
$\mathbb{P}(\mathcal{V})\leq\alpha.$
Proof.
For each step $t$, consider the event
Observe that
We are now ready to present the proof of Theorem 3 using Lemma 6.
Theorem 3.
The proof of this result follows from observing that

### v2

@@ 6 Conclusion
[…]
Before presenting additional results on clustering quality, we first provide the definitions of the clustering quality metrics used in our study.
  1. (i)
  2. (ii)
    $\text{ARI}=\dfrac{\sum_{k,k^{\prime}}\binom{n_{k,k^{\prime}}}{2}-\left[\sum_{k}\binom{a_{k}}{2}\sum_{k^{\prime}}\binom{b_{k^{\prime}}}{2}\right]\Big/\binom{n}{2}}{\dfrac{1}{2}\left[\sum_{k}\binom{a_{k}}{2}+\sum_{k^{\prime}}\binom{b_{k^{\prime}}}{2}\right]-\left[\sum_{k}\binom{a_{k}}{2}\sum_{k^{\prime}}\binom{b_{k^{\prime}}}{2}\right]\Big/\binom{n}{2}}$,
[…]
Continuing the simulations in Section 5.1, we now present additional results assessing the clustering performance of Algorithm 1. Specifically, in addition to comparing performance across different randomization level, we compare the performance of the randomized hierarchical clustering algorithm under varying levels of signal strength $\delta$, while fixing the randomization level at $\tau^{*}=0.10$ (i.e., implementing RC(3)). Figure 7 shows that the randomized hierarchical clustering algorithm with $\tau^{*}=0.10$ achieves performance very similar to the deterministic version across a wide range of signal strengths $\delta$. When the clusters are well separated (larger $\delta$), the randomized method attains nearly perfect recovery of the true clustering, as reflected by low WCSS/TSS ratios and high ARI values. Overall, these results suggest that a small amount of randomization does not degrade clustering quality, even as the signal strength varies. The benefits of the added randomization, however, become evident in the next sections, where we demonstrate the performance of our testing procedures for validating the clusters discovered by Algorithm 1.
[…]
We begin this section by presenting additional experiment results for the experiments in Section 5.2.
Following the setting described in Section 5.2, generate data under the null hypothesis and partition data into $K=3$ clusters. Figure 8 shows that the proposed method produces uniform p-values under the null hypothesis and successfully controls the Type I error rate.
To evaluate power of the proposed p-value using $K=3$, we use the data generated from three equidistant clusters
Figure 9 illustrates that the proposed method consistently achieves greater power compared to existing selective inference procedures.
Figure 9: Empirical power curves as a function of effect size for the proposed randomized method with $\tau^{*}=0.10$, compared with the two selective inference approaches under varying choice of linkage functions (complete, single, average and minimax) and true number of clusters $K=3$.
Next, we present the comparison of the proposed p-value validity with existing selective inference methods. Using the same data-generating model as in Section 5.2 with $n=30$, $p=10$, $\sigma=1$, we perform our clustering algorithm until the data are partitioned into $K=2$ and $K=3$ clusters, respectively, under four linkage criteria: complete, average, single, and minimax. For each setting, we perform $2000$ repetitions and evaluate the ECDF of the resulting p-values to assess validity under the null. The software implementations of the methods proposed by [Gao et al., 2024] and [Yun and Barber, 2023] do not support testing clusters obtained using the minimax linkage criterion. Figure 10 shows that, across all settings, the ECDFs of the p-values from RC(3) align closely with the reference uniform distribution, indicating validity. As expected, the results are similar to the two existing selective inferential methods whenever these approaches are available for a given linkage criterion.
Figure 10: Comparison of the ECDFs of p-values obtained under the null hypothesis. The proposed method with $\tau^{*}=0.10$ is compared against [Gao et al., 2024] and [Yun and Barber, 2023], across different linkage criteria and cluster numbers $K\in\{2,3\}$.
[…]
We further evaluate the performance of Algorithm 2 under varying true numbers of clusters $K^{*}\in\{1,2,\dots,10\}$. For each value of $K^{*}$, we generate $100$ independent datasets with $n=200$ observations and $p=2$ dimensions. The $K^{*}$ cluster centers are placed evenly around a circle of radius $\delta=6$, and each data point is sampled from a spherical Gaussian distribution centered at its corresponding cluster. Cluster sizes are distributed as evenly as possible across groups. We use the same procedure as in the experiments in Section 5.3 to generate the $\alpha$-sequence, and set $n_{\min}=10$ and $n^{*}=40$.